# From MRI to SARI: Parametric survey of non-axisymmetric accretion disk instabilities

### v1

@@ 5.4.1 Instability criterion
Much of the SARI behaviour can be explained by the instability criterion (10). Contrary to the $m=0$ MRI, the toroidal field plays a significant role in the instability of non-axisymmetric modes. An important consequence is that the perturbations can couple to the toroidal field: for any given value of $k$, there exists a negative $m$ that makes $\omega_{A}$ small enough to satisfy the criterion. A second consequence is that for certain wavenumber combinations, $\omega_{A}(r_{*})=0$ locally, hence violating the lower end of the instability criterion. Both of these possibilities are completely absent for the MRI.
@@
A direct consequence of the first observation is the shift in cut-off vertical wavenumber with $m$ as found in Fig. 4. From Eq. (10), it is clear that the cut-off vertical wavenumber shifts to larger values compared to the axisymmetric MRI for $m<0$ and to smaller values for $m>0$. For a sizeable poloidal field, this effect is small: for Case (a), which was not shown, all cut-offs are in the interval $[150,175]$ for $|m|\leq 10$. For a dominant poloidal field, the effect is more noticeable: according to the analytical predictions, $k_{\text{max}}$ and $k_{\text{cut-off}}$ shift to large values when $m$ becomes arbitrarily negative. Furthermore, the wavenumber $k$ where $\omega_{A}$ vanishes locally shifts to larger values as $m$ becomes increasingly negative. For increasingly positive $m$, the range of allowed $k$ decreases due to stabilisation by the increasing magnetic tension.
@@
Figure 9: Alternative view of Fig. 4, where the inner SARI growth rates now depend on $\omega_{A}$ varying through $k$. Instability is perfectly governed by the criterion $0<\omega_{A}^{2}\lesssim 3\Omega_{1}^{2}$. Most growth rates peak around $\omega_{A}\approx\sqrt{15/16}\ \Omega_{1}$.
@@
The validity of the instability criterion is made strikingly clear when we plot the growth rates as a function of $\omega_{A}(r_{1})$, as in Fig. 9. It contains the same results as Fig. 4 where $k$ is varied, without the outer SARIs. Stability is indeed approximately limited to the region where $\omega_{A}(r_{1})<\sqrt{3}\Omega_{1}$ and it is clear how the $m<0$ SARIs have access to the whole unstable range, and are stabilised when the Alfvén frequency moves through zero. Note how the $m>0$ SARIs cannot access the whole range of unstable Alfvén frequencies. The maximum growth rates are indeed approximately reached for $\omega_{A}(r_{1})=\sqrt{15/16}\ \Omega_{1}$, as predicted by Eq. (11), but this value shifts depending on $m$. The theoretical maximum growth of $0.75\Omega_{1}$ is, however, not attained. The slight differences for $\omega_{A\text{max}}$ and the maximum growth rate depending on $m$ is partly explained by the fact that these inner SARI growth rates are not evaluated at the exact Doppler radius where their eigenfunctions peak, but at the inner radius instead. Finally, note that the equilibrium for Fig. 9 has $v_{A1}\sim\epsilon\ll 1\sim r_{1}\Omega_{1}$ and hence is in the super-Alfvénic regime. This is not in contradiction with Fig. 9: indeed, although instability extends to wavenumbers $k$ where $\omega_{A1}>\Omega_{1}$, the rotation frequency in the local Doppler frame $\omega_{A1}<m\Omega_{1}$ is still super-Alfvénic.
@@
The outer SARI growth rates (not shown) obey a similar dependency on $\omega_{A}(r_{2})$ when scaled by $\Omega(r_{2})$. Note also that the condition $\omega_{A}^{2}(r_{2})<3\Omega^{2}(r_{2})$ for outer SARI instability is more restrictive than that for the inner SARIs, as the right hand side of the inequality decreases faster than the left hand side. This explains why outer SARI modes are stabilised more easily compared to inner SARIs. We also noted in the previous Sections that outer SARIs have growth rates three times smaller than inner SARIs. When the outer SARIs are, however, scaled by the local rotation frequency $\approx\Omega(r_{2})$ instead of $\Omega_{1}$, this discrepancy disappears. Indeed, $\Omega(r_{2})=\Omega_{1}r_{2}^{-3/2}\approx\Omega_{1}/3$ if $\delta=1$. Relative to the local frame of reference, the outer modes are hence equally fast-growing over one rotation period, with maximum growth rates up to $0.7\Omega$ in our numerical results. The fact that the outer SARIs are more easily stabilised than the inner modes, gives two possibilities for the evolution of the entire spectrum when an equilibrium parameter is varied: (1) the quasi-continuum first collapses to a point as the overlap between the Alfvén continua decreases, where it might detach from the real axis. The inner and outer branches then merge into one branch of discrete modes, or the outer branch is stabilised before that happens; or (2) the outer branch is stabilised before the quasi-continuum disappears, and a quasi-continuum still exists as long as the overlap of the continua is large enough. In either case, the remaining branch of inner modes is reminiscent of the unstable sequence of MRI modes, but it contains an infinite number of modes clustering to an internal edge of one of the continua $\Omega_{A}^{\pm}$ (Goedbloed & Keppens, 2022). It, too, eventually disappears to the real axis when it is stabilised.
@@
To compare the predictions to the actual values of $k_{\text{cut-off}}$ of Figs. 48, we evaluate Eq. (10) at the local Doppler radius $r_{*}$ where $\sigma=\Omega_{0}(r_{*})$, and MRI growth rates at the inner boundary $r_{1}$. These predictions are reasonably accurate when compared to the actual growth rates of Fig. 4: e.g. the $m=-3$ SARI is predicted to have $k_{\text{cut-off}}=411$ while in reality $k_{\text{cut-off}}=400$, and the value of $k$ where the inner mode is locally stabilised is predicted to be $k_{\text{stab}}=33$ while in reality it happens as $k_{\text{stab}}=30$. When keeping $k$ fixed in Figs. 5 and 7, it is the orientation or strength of the magnetic field that will determine whether Eq. (10) is satisfied. The Alfvén frequency in terms of the equilibrium parameters is given by $\omega_{A}^{2}(r)=\left(m\mu_{1}/r+k\right)\epsilon\left(1+\mu_{1}^{2}\right)^{-1/2}r^{-1/2}$. Hence, the instability criterion predicts that, once the field gets stronger, large values of $k$ are only allowed when the toroidal component is significant; this is exactly the difference between Fig. 5 (a) and (b), which both have $k=50$. The instability criterion again explains the ordering of the cut-off vertical field obtained for Case (b): if $m<0$, $B_{\theta}$ can become smaller compared to $B_{z}$, and vice-versa for $m>0$. For Case (a), there is no cut-off as $k$ is below the critical value $k_{\text{cut-off}}$ for a purely poloidal field. For a fixed orientation and wavenumber, instability is only possible in a range of field strengths where the upper limit of Eq. (10) is satisfied. The ordering of $m<0$ and $m>0$ mode critical field strengths in Figs. 7 (a) and (b) is again explained from this criterion, except for large $v_{A1}$ in Case (b). Interestingly, Eq. (10) leaves the option for MRI/SARI in a range of wavenumbers to be ‘switched off’ at intermediate pitch angles $\varphi$, since $\omega_{A}$ as a function of $\varphi$ might first increase above the critical threshold of Eq. (10) and then again decrease. This implies that some high-$m$ co-rotating SARIs are stabilised for e.g. $\varphi\in[5^{\circ},25^{\circ}]$, which means that they are unstable for an almost purely vertical field or when a significant azimuthal field is present. Finally, a visualisation of $\omega_{A}$ as $\mu_{1}$ or $\epsilon$ are varied gives almost the same results as in Fig. 9.
@@ 5.4.2 Growth rate estimates
We now investigate the upper bound to the SARI growth rates by evaluating Eq. (9) at the local Doppler corotation radius where $\sigma=\Omega_{0}(r_{*})$, which is also where the eigenfunctions peak. Expression (9) for the maximal growth rate varies with $r$ and is usually radially decreasing for our accretion disk equilibria. This implies that inner SARIs (localised around $r_{1}$) indeed have larger growth rates than outer SARIs (localised around $r_{2}$), as was found in most of our spectra and parameter scans. There are, however, specific instances where $\nu_{\text{max}}$ is non-decreasing and the outer SARI can have larger growth rates. Recall for example the cases with vanishing $\omega_{A}(r_{1})$ above, or the $m=-3$ SARI in Fig. 8. In the second case, $\nu_{\text{max}}$ increases from its value at $r_{1}$ up to a global maximum, after which it again decreases to its value in $r_{2}$, which is larger than that in $r_{1}$. However, there is no mode corresponding to the maximum of $\nu_{\text{max}}$ with a growth rate larger than that of the outer mode, so the outer SARI is the most unstable mode. This also points at a second observation: the predicted upper bound for growth rates is never strict. It usually overestimates the growth rates, both for discrete and quasi-continuum modes, but one can nevertheless derive the shape of the quasi-continuum upper edge from it, as in Fig. 10. A possible reason for this discrepancy is that the derivative of the eigenfunctions, $\chi^{\prime}$, was not taken into account in the derivation of (9), as noted in Goedbloed & Keppens (2022). We will now use Eqs. (9) and (10) to explain several observations made earlier.
@@
To fully confirm the findings of Fig. 6 that a purely toroidal field allows for arbitrarily large unstable wavenumbers $k$, we extend the numerical results with predictions using Eq. (9). Indeed, as $\varphi\rightarrow 90^{\circ}$, the values of $k_{\text{max}}$ and of $k_{\text{cut-off}}$ increase dramatically: the growth rates for $\varphi=89.99^{\circ}$ are predicted to peak at $k=2\times 10^{5}$, and for $\varphi=88.999^{\circ}$ at $k=1.5\times 10^{6}$; the most unstable modes can become arbitrarily localised in the vertical direction. Here, the distinction with the $m=0$ MRI becomes clear: we performed the same Legolas runs as in Fig. 6 but now for the MRI, extended with analytical predictions for very large $k$. Unsurprisingly, the $m=0$ MRI also shows increased localisation as $\varphi\rightarrow 90^{\circ}$ because the poloidal component gets weaker and hence larger $k$ are allowed. However, whereas the SARI retains finite growth rates for small $k$ as $\varphi\rightarrow 90^{\circ}$, the MRI is stabilised in the limit: $k_{\text{max}}$ moves to arbitrarily large values, but for values of $k\leq 1000$, the growth rates become arbitrarily small. In contrast to the SARI modes, the $m=0$ MRI has no way to couple to the toroidal field, and once the poloidal component becomes negligible, it is stabilised.
@@
Expression (9) for $\nu_{\text{max}}$ also explains why ±mplus-or-$\pm m$ SARIs have coalescing growth rates for $\varphi\rightarrow 0^{\circ}$ or $\varphi\rightarrow 90^{\circ}$ in Fig. 5 for both Case (a) and Case (b). Indeed, for a purely toroidal field, $\nu_{\text{max}}$ only depends on $m^{2}$ and for a purely poloidal field, $m$ completely disappears from the expression (but evaluating $\nu_{\text{max}}$ at the proper resonance location $r_{*}$ for each mode does produce the observed dependence on $m$). Physically, a purely vertical or toroidal field induces no preferential (co- or counter-rotating) direction for waves to grow. As mentioned in Sect. 5.3, Fig. 7 (b) shows a peculiar ordering of the critical field strengths $v_{A1\text{cut-off}}$, which are decreasing as $m$ increases from $m=-3$ to $10$, but the $m=-10$ SARI has a smaller range of unstable field strengths. From Eq. (9), we find the following behaviour of the instability as $m$ varies: for $m>0$, the range of unstable field strengths decreases and disappears in the limit $m\rightarrow\infty$, since then $\omega_{A}$ becomes large. For $m<0$, the critical field strength first increases to a maximum, then decreases and eventually vanishes too in the limit $m\rightarrow-\infty$. When the field has a dominant toroidal component, this switchback occurs for lower $m$ than when the field is dominantly poloidal.
@@
We conclude that, in general, Eq. (9) correctly predicts the general trend of growth rates in weakly magnetised, thin disks, the parameter regime for which it was derived. Outside of this regime, it breaks down: when the assumption $k^{2}r^{2}\gg m^{2}$ is violated in Fig. 4 all growth rates nearly vanish as $k\rightarrow 0$ but the predicted growth rates remain finite; for $\beta<1$ in Fig. 8, growth rates are not at all correctly predicted, and the $m=\pm 10$ SARIs behave in a distinct way. It is clear that in the compressible regime, one has to resort to numerical means as done in this work. Equation (9) does not replicate the decreasing maximum growth rates as $|m|$ increases, which was observed above in most figures.

### v2

@@ 5.4.1 Instability criterion
Much of the SARI behaviour can be explained by the instability criterion (10). Contrary to the $m=0$ MRI, the toroidal field plays a significant role in the instability of non-axisymmetric modes. An important consequence is that the perturbations can couple to the toroidal field: for any given value of $k$, there exists a negative $m$ that makes $\omega_{A}$ low enough to satisfy the criterion. A second consequence is that for certain wavenumber combinations, $\omega_{A}(r_{*})=0$ locally, hence violating the lower end of the instability criterion. Both of these possibilities are completely absent for the MRI.
@@
A direct consequence of the first observation is a shift in cut-off vertical wavenumber with $m$ as found in Fig. 4. From Eq. (10), it is clear that the cut-off vertical wavenumber shifts to larger values compared to the axisymmetric MRI for $m<0$ and to smaller values for $m>0$. For a sizeable poloidal field, this effect is small: for Case (a), which is not shown, all cut-offs are in the interval $[150,175]$ for $|m|\leq 10$. For a dominant poloidal field, the effect is more noticeable: according to the analytical predictions, $k_{\text{max}}$ and $k_{\text{cut-off}}$ shift to large values when $m$ becomes arbitrarily negative. Furthermore, the wavenumber $k$ where $\omega_{A}$ vanishes locally shifts to larger values as $m$ becomes increasingly negative. For increasingly positive $m$, the range of allowed $k$ decreases due to stabilisation by the increasing magnetic tension.
@@
Figure 9: Alternative view of Fig. 4, where the inner SARI growth rates now depend on $\omega_{A}$ varying through $k$. Instability is perfectly governed by the criterion $0<\omega_{A}^{2}\lesssim 3\Omega_{1}^{2}$. Most growth rates peak at around $\omega_{A}\approx\sqrt{15/16}\ \Omega_{1}$.
@@
The validity of the instability criterion is made strikingly clear when we plot the growth rates as a function of $\omega_{A}(r_{1})$, as in Fig. 9. It contains the same results of Fig. 4 where $k$ is varied, but without the outer SARIs. Stability is indeed approximately limited to the region where $\omega_{A}(r_{1})<\sqrt{3}\Omega_{1}$ and it is clear how the $m<0$ SARIs have access to the whole unstable range, and are stabilised when the Alfvén frequency moves through zero. Note how the $m>0$ SARIs cannot access the whole range of unstable Alfvén frequencies. The maximum growth rates are indeed approximately reached for $\omega_{A}(r_{1})=\sqrt{15/16}\ \Omega_{1}$, as predicted by Eq. (11), but this value shifts depending on $m$. The theoretical maximum growth of $0.75\Omega_{1}$ is, however, not attained. The slight differences for $\omega_{A\text{max}}$ and the maximum growth rate depending on $m$ are partly explained by the fact that these inner SARI growth rates are evaluated not at the exact Doppler radius where their eigenfunctions peak, but at the inner radius instead. Finally, note that the equilibrium for Fig. 9 has $v_{A1}\sim\epsilon\ll 1\sim r_{1}\Omega_{1}$ and hence is in the super-Alfvénic regime. This is not in contradiction with Fig. 9: indeed, although instability extends to wavenumbers $k$ where $\omega_{A1}>\Omega_{1}$, the rotation frequency in the local Doppler frame $\omega_{A1}<m\Omega_{1}$ is still super-Alfvénic.
@@
The outer SARI growth rates (not shown) obey a similar dependency on $\omega_{A}(r_{2})$ when scaled by $\Omega(r_{2})$. Note also that the condition $\omega_{A}^{2}(r_{2})<3\Omega^{2}(r_{2})$ for outer SARI instability is more restrictive than that for the inner SARIs, as the right-hand side of the inequality decreases faster than the left-hand side. This explains why outer SARI modes are stabilised more easily compared to inner SARIs. We also noted in the previous Sections that outer SARIs have growth rates three times lower than those of inner SARIs. When the outer SARIs are, however, scaled by the local rotation frequency $\approx\Omega(r_{2})$ instead of $\Omega_{1}$, this discrepancy disappears. Indeed, $\Omega(r_{2})=\Omega_{1}r_{2}^{-3/2}\approx\Omega_{1}/3$ if $\delta=1$. Relative to the local frame of reference, the outer modes are hence equally fast-growing over one rotation period, with maximum growth rates up to $0.7\Omega$ in our numerical results. The fact that the outer SARIs are more easily stabilised than the inner modes gives two possibilities for the evolution of the entire spectrum when an equilibrium parameter is varied: (1) the quasi-continuum first collapses to a point as the overlap between the Alfvén continua decreases, where it might detach from the real axis. The inner and outer branches then merge into one branch of discrete modes, or the outer branch is stabilised before that happens; or (2) the outer branch is stabilised before the quasi-continuum disappears, and the quasi-continuum still exists as long as the overlap of the continua is large enough. In either case, the remaining branch of inner modes is reminiscent of the unstable sequence of MRI modes, but it contains an infinite number of modes clustering to an internal edge of one of the continua $\Omega_{A}^{\pm}$ (Goedbloed & Keppens, 2022). It, too, eventually disappears to the real axis when it is stabilised.
@@
To compare the predictions to the actual values of $k_{\text{cut-off}}$ in Figs. 48, we evaluate Eq. (10) at the local Doppler radius $r_{*}$ where $\sigma=\Omega_{0}(r_{*})$, and the MRI growth rates at the inner boundary $r_{1}$. These predictions are reasonably accurate when compared to the actual growth rates of Fig. 4: e.g. the $m=-3$ SARI is predicted to have $k_{\text{cut-off}}=411$ while in reality $k_{\text{cut-off}}=400$, and the value of $k$ where the inner mode is locally stabilised is predicted to be $k_{\text{stab}}=33$ while in reality it happens as $k_{\text{stab}}=30$. When keeping $k$ fixed in Figs. 5 and 7, it is the orientation or strength of the magnetic field that determines whether Eq. (10) is satisfied. The Alfvén frequency in terms of the equilibrium parameters is given by $\omega_{A}^{2}(r)=\left(m\mu_{1}/r+k\right)\epsilon\left(1+\mu_{1}^{2}\right)^{-1/2}r^{-1/2}$. Hence, the instability criterion predicts that, once the field gets stronger, large values of $k$ are only allowed when the toroidal component is significant; this is exactly the difference between Fig. 5 (a) and (b), which both have $k=50$. The instability criterion again explains the ordering of the cut-off vertical field obtained for Case (b): if $m<0$, $B_{\theta}$ can become smaller compared to $B_{z}$, and vice-versa for $m>0$. For Case (a), there is no cut-off as $k$ is below the critical value $k_{\text{cut-off}}$ for a purely poloidal field. For a fixed orientation and wavenumber, instability is only possible in a range of field strengths where the upper limit of Eq. (10) is satisfied. The ordering of $m<0$ and $m>0$ mode critical field strengths in Figs. 7 (a) and (b) is again explained from this criterion, except for the high $v_{A1}$ in Case (b). Interestingly, Eq. (10) leaves the option for MRI/SARI in a range of wavenumbers to be ‘switched off’ at intermediate pitch angles $\varphi$, since $\omega_{A}$ as a function of $\varphi$ might first increase above the critical threshold of Eq. (10) and then again decrease. This implies that some high-$m$ co-rotating SARIs are stabilised for $\varphi\in[5^{\circ},25^{\circ}]$, for example, which means that they are unstable for an almost purely vertical field or when a significant azimuthal field is present. Finally, a visualisation of $\omega_{A}$ as $\mu_{1}$ or $\epsilon$ is varied gives almost the same results as those in Fig. 9.
@@ 5.4.2 Growth rate estimates
We now investigate the upper bound to the SARI growth rates by evaluating Eq. (9) at the local Doppler corotation radius, where $\sigma=\Omega_{0}(r_{*})$, which is also where the eigenfunctions peak. Expression (9) for the maximal growth rate varies with $r$ and is usually radially decreasing for our accretion disk equilibria. This implies that inner SARIs (localised around $r_{1}$) indeed have higher growth rates than outer SARIs (localised around $r_{2}$), as was found in most of our spectra and parameter scans. There are, however, specific instances where $\nu_{\text{max}}$ is non-decreasing and the outer SARI can have higher growth rates. Recall for example the cases with vanishing $\omega_{A}(r_{1})$ above, or the $m=-3$ SARI in Fig. 8. In the second case, $\nu_{\text{max}}$ increases from its value at $r_{1}$ up to a global maximum, after which it again decreases to its value in $r_{2}$, which is larger than that in $r_{1}$. However, there is no mode corresponding to the maximum of $\nu_{\text{max}}$ with a growth rate higher than that of the outer mode, so the outer SARI is the most unstable mode. This also points to a second observation: the predicted upper bound for growth rates is never strict. It usually overestimates the growth rates, for both discrete and quasi-continuum modes, but one can nevertheless derive the shape of the quasi-continuum upper edge from it, as in Fig. 10. A possible reason for this discrepancy is that the derivative of the eigenfunctions, $\chi^{\prime}$, was not taken into account in the derivation of (9), as noted in Goedbloed & Keppens (2022). We will now use Eqs. (9) and (10) to explain several observations made earlier.
@@
To fully confirm the finding of Fig. 6 that a purely toroidal field allows for arbitrarily large unstable wavenumbers $k$, we extend the numerical results with predictions using Eq. (9). Indeed, as $\varphi\rightarrow 90^{\circ}$, the values of $k_{\text{max}}$ and of $k_{\text{cut-off}}$ increase dramatically: the growth rates for $\varphi=89.99^{\circ}$ are predicted to peak at $k=2\times 10^{5}$, and for $\varphi=88.999^{\circ}$ at $k=1.5\times 10^{6}$; the most unstable modes can become arbitrarily localised in the vertical direction. Here, the distinction with the $m=0$ MRI becomes clear: we perform the same Legolas runs as in Fig. 6 but now for the MRI, extended with analytical predictions for very large $k$. Unsurprisingly, the $m=0$ MRI also shows increased localisation as $\varphi\rightarrow 90^{\circ}$ because the poloidal component gets weaker and hence larger $k$ are allowed. However, whereas the SARI retains finite growth rates for small $k$ as $\varphi\rightarrow 90^{\circ}$, the MRI is stabilised in the limit: $k_{\text{max}}$ moves to arbitrarily large values, but for values of $k\leq 1000$, the growth rates become arbitrarily low. In contrast to the SARI modes, the $m=0$ MRI has no way to couple to the toroidal field, and once the poloidal component becomes negligible, it is stabilised.
@@
Expression (9) for $\nu_{\text{max}}$ also explains why ±mplus-or-$\pm m$ SARIs have coalescing growth rates for $\varphi\rightarrow 0^{\circ}$ or $\varphi\rightarrow 90^{\circ}$ in Fig. 5 for both Case (a) and Case (b). Indeed, for a purely toroidal field, $\nu_{\text{max}}$ only depends on $m^{2}$ and for a purely poloidal field, $m$ completely disappears from the expression (but evaluating $\nu_{\text{max}}$ at the proper resonance location $r_{*}$ for each mode does produce the observed dependence on $m$). Physically, a purely vertical or toroidal field induces no preferential (co- or counter-rotating) direction for waves to grow in. As mentioned in Sect. 5.3, Fig. 7 (b) shows a peculiar ordering of the critical field strengths $v_{A1\text{cut-off}}$, which decrease as $m$ increases from $m=-3$ to $10$, but the $m=-10$ SARI has a smaller range of unstable field strengths. From Eq. (9), we find the following behaviour of the instability as $m$ varies: for $m>0$, the range of unstable field strengths decreases and disappears in the limit $m\rightarrow\infty$, since then $\omega_{A}$ becomes high. For $m<0$, the critical field strength first increases to a maximum, then decreases and eventually vanishes too in the limit $m\rightarrow-\infty$. When the field has a dominant toroidal component, this switchback occurs for lower $m$ than when the field is dominantly poloidal.
@@
We conclude that, in general, Eq. (9) correctly predicts the general trend of growth rates in weakly magnetised, thin disks, the parameter regime for which it was derived. Outside of this regime, it breaks down: when the assumption $k^{2}r^{2}\gg m^{2}$ is violated in Fig. 4 all growth rates nearly vanish as $k\rightarrow 0$ but the predicted growth rates remain finite; for $\beta<1$ in Fig. 8, growth rates are not at all correctly predicted, and the $m=\pm 10$ SARIs behave in a distinct way. It is clear that in the compressible regime, one has to resort to numerical means as done in this work. Equation (9) does not replicate the decreasing maximum growth rates as $|m|$ increases, which is observed above in most figures.